Hello my name Kenny Adams and I am having a problem that really dosent make since to me the problem is x+y=180,p+q= 180 and p = x Show that y =a I found this problem on the USA TEST PREP for the HSAP and i could not no where to start I tried to work it out but every time I try it is unsuccseful I need help could u at least tell me how to start because in my 10 years of school none of my teachers have ever shown me a problem like this and it is required that I know these kind of problems to past the test I would really appreciate if u could get me stared Thank You
Found 2 solutions by stanbon, bucky:
Answer by stanbon(48526)

Something is missing in your post or in your copy of the
problem.
You say nothing about how "a" is related to anything.
If x+y=180
and x=p
then you can substitute and get p+y=180
But that doesn't tell you anything about "a"
because nothing has been said about "a".
The problem could just as well say: Prove y= z or y=w.
Look again at the problem. What does it say about "a"
before it says: "Show that y=a".
Cheers,
Stan H.

Answer by bucky(2097)

Hi Kenny
.
I'll try to help.
.
Think about this ... You are given two equations
.
x + y = 180 and
p + q = 180
.
Notice that if the right sides of these two equations are equal (and they are both 180),
then the left sides have to be equal to each other too since they both have to equal 180.
Because of this we can set the left sides equal like this:
.
x + y = p + q
.
Then the problem says that p is equal to x. So in this equation we can plug x in wherever
we see p. Do that and the equation becomes:
.
x + y = x + q
.
Since we now have an x on both sides of the equation, we can get rid of the x terms by
subtracting x from both sides. (Remember, we can add to an equation, subtract from an
equation, multiply all the terms in an equation, or divide all the terms in an
equation
as long as we do the same thing to BOTH sides of the equation.) This time we are going
to subtract x, so we need to subtract x from BOTH sides of the equation.
.
When we subtract (take away) an x from both sides the x terms on each side are gone, and
the equation we are left with is just:
.
y = q
.
Your problem says you are to show that y = a, but I think that's a keyboarding mistake and
you meant to type y = q (q is the keyboard letter right above the letter "a" so typing "a"
instead of q would be easy to do).
.
Hope this helps you to understand the problem and that you do well on the exam. Keep working
at it!
.
Bucky